# MADAR: Efficient Continual Learning for Malware Analysis with Diversity-Aware Replay

### v1

@@ VIII Conclusion
In this paper, we propose MADAR, a framework for diversity-aware replay in continual learning specially designed for the challenging setting of malware classification. Our comprehensive evaluation across Domain-IL, Class-IL, and Task-IL scenarios against Windows executable (EMBER) and Android application (AZ) datasets demonstrates that diversity-aware sampling is helpful for effective CL in malware classification. As malware and goodware continue to evolve, these insights steer continual learning towards strategic, resource-efficient methods, ensuring model effectiveness amid the constantly shifting landscape of cybersecurity threats.

### v2

@@ VIII Conclusion
In this paper, we propose MADAR, a framework for distribution-aware replay in continual learning specially designed for the challenging setting of malware classification. Our comprehensive evaluation across Domain-IL, Class-IL, and Task-IL scenarios against Windows executable (EMBER) and Android application (AZ) datasets demonstrates that distribution-aware sampling is helpful for effective CL in malware classification. As malware and goodware continue to evolve, these insights steer continual learning towards strategic, resource-efficient methods, ensuring model effectiveness amid the constantly shifting landscape of cybersecurity threats.
Acknowledgements
This research was funded in part by the National Science Foundation under Grant no. 2422241. The authors acknowledge Research Computing [42] at the Rochester Institute of Technology for providing computational resources and support.